# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

### v1

@@ 4.3.1 Artificial data
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.000$ $\bf{0.927}$ $0.073$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.243)$ $(0.243)$ $(0.000)$ $(0.000)$
$\alpha=0.2$ $0.000$ $0.012$ $\bf{0.966}$ $0.022$ $0.000$ $0.000$
$(0.000)$ $(0.084)$ $(0.111)$ $(0.073)$ $(0.001)$ $(0.000)$
$\alpha=10^{-1}$ $0.000$ $0.020$ $\bf{0.950}$ $0.027$ $0.002$ $0.000$
$(0.000)$ $(0.141)$ $(0.178)$ $(0.106)$ $(0.009)$ $(0.002)$
$\alpha=10^{-2}$ $0.000$ $0.036$ $\bf{0.936}$ $0.023$ $0.004$ $0.001$
$(0.000)$ $(0.179)$ $(0.184)$ $(0.035)$ $(0.008)$ $(0.004)$
$\alpha=10^{-3}$ $0.000$ $0.028$ $\bf{0.938}$ $0.023$ $0.007$ $0.004$
$(0.000)$ $(0.143)$ $(0.180)$ $(0.044)$ $(0.021)$ $(0.017)$
@@
$\gamma=1$ $0.000$ $0.020$ $\bf{0.930}$ $0.049$ $0.002$ $0.000$
$(0.000)$ $(0.139)$ $(0.135)$ $(0.028)$ $(0.002)$ $(0.000)$
$\gamma=0.2$ $0.000$ $0.010$ $\bf{0.480}$ $0.307$ $0.133$ $0.070$
$(0.000)$ $(0.069)$ $(0.085)$ $(0.036)$ $(0.035)$ $(0.038)$
$\gamma=10^{-1}$ $0.000$ $0.035$ $\bf{0.300}$ $0.288$ $0.186$ $0.191$
$(0.000)$ $(0.104)$ $(0.059)$ $(0.042)$ $(0.040)$ $(0.070)$
$\gamma=10^{-2}$ $0.000$ $0.010$ $0.091$ $0.156$ $\bf{0.175}$ $0.566$
$(0.000)$ $(0.041)$ $(0.036)$ $(0.014)$ $(0.015)$ $(0.092)$
$\gamma=10^{-3}$ $0.000$ $0.006$ $0.069$ $0.131$ $\bf{0.160}$ $0.634$
$(0.000)$ $(0.028)$ $(0.023)$ $(0.017)$ $(0.012)$ $(0.085)$
From table 8, MFM-Inv-Ga has a much higher posterior probability at $M_{\text{true}}$ for all shape parameters than MFM-Ga. As a result, in the context of community detection, MFM-Inv-Ga also achieves better clustering and community estimation than MFM-Ga.
We also compared the computation time with Geng’s method (Geng et al., 2019), denoted MFM-Geng, when the shape parameters are $1.0$ and $10^{-3}$. The result in Table 9 indicates that MFM-Inv-Ga is on average more than twice as efficient as MFM-Geng in terms of computation time. The reason is why our algorithm does not update the label variable based on the restaurant process and MFM-Inv-Ga has a structure that is unlikely to produce empty components.
The MFM-Geng can estimate the number of clusters $k$, but cannot directly estimate the number of communities $M$. Furthermore, it is difficult in the MFM-Geng to estimate the hyperparameter of $q_{M}$, because it is necessary to perform complex series calculations, including its parameter. In summary, MFM-Inv-Ga (and MFM-Ga) are superior to MFM-Geng in terms of computation time, direct estimation of $M$, and estimation of an essential parameter.
Table 9: Comparison of CPU times in seconds (standard deviations are shown in parentheses) averaged over $50$ replications.
MFM-Inv-Ga MFM-Geng
$\alpha=\gamma=1$ $38.793$ $87.519$
$(2.894)$ $(3.719)$
$\alpha=\gamma=10^{-3}$ $38.361$ $83.423$
$(4.431)$ $(4.673)$
4.3.2 Dolphins social network data
The dolphins social network data is often used as a benchmark. Data can be obtained in http://www-personal.umich.edu/mejn/netdata/. The data is constructed as an undirected graph and expresses a small-scale animal social network with $64$ bottlenose dolphins off Doubtful Sound, New Zealand. Each node represents a dolphin, and an edge is drawn if two dolphins appear to be closely related to each other. In previous studies, it is well-known that the network has two communities. The details of the data are found in Lusseau et al. (2003).
As before, we compare the posterior distribution of the number of communities between MFM-Inv-Ga and MFM-Ga, and create a co-clustering matrix of MFM-Inv-Ga as quantifying uncertainty of clustering. In this analysis, we set $M-1\sim\text{Poisson}(1)$ and $a_{Q}=b_{Q}=3.0$.
Table 10: Posterior probabilities of the number of communities $M$ for dolphins social network data
MFM-Inv-Ga
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M\geq 6$
$\alpha=1$ $0.000$ $0.997$ $0.003$ $0.000$ $0.000$ $0.000$
$\alpha=0.2$ $0.000$ $0.972$ $0.025$ $0.002$ $0.001$ $0.000$
$\alpha=10^{-1}$ $0.000$ $0.961$ $0.034$ $0.004$ $0.000$ $0.000$
$\alpha=10^{-2}$ $0.000$ $0.941$ $0.048$ $0.010$ $0.001$ $0.000$
$\alpha=10^{-3}$ $0.000$ $0.854$ $0.136$ $0.009$ $0.002$ $0.000$
MFM-Ga
$\gamma=1$ $0.000$ $0.949$ $0.049$ $0.002$ $0.000$ $0.000$
$\gamma=0.2$ $0.000$ $0.605$ $0.307$ $0.074$ $0.012$ $0.001$
$\gamma=10^{-1}$ $0.000$ $0.514$ $0.340$ $0.117$ $0.025$ $0.004$
$\gamma=10^{-2}$ $0.000$ $0.619$ $0.303$ $0.069$ $0.000$ $0.000$
$\gamma=10^{-3}$ $0.370$ $0.400$ $0.172$ $0.048$ $0.007$ $0.001$
From Table 10, the MFM-Inv-Ga is able to successfully estimate the number of communities $M=2$ regardless of the values of the shape parameter, while the MFM-Ga is not. We report a co-cluster matrix of MFM-Inv-Ga for $\alpha=1.0$ and MFM-Geng for $\gamma=1.0$, and the results are shown in Figure 5. The result of MFM-Inv-Ga is almost identical to the results reported in Geng et al. (2019). Furthermore, we confirmed that changing the value of $\alpha$ does not change the co-cluster matrices and the clustering solution based on MAP estimation. This shows the efficiency of the proposed method.
Figure 5: Co-clustering matrix with the MFM-Inv-Ga with $\alpha=1.0$ (left) and the MFM-Geng with $\gamma=1.0$ (right) for dolphins social network data.
5 Concluding remarks
We proposed a mixture of finite mixtures (MFM) model based on the normalized inverse Gaussian distribution and constructed an efficient posterior sampling algorithm based on Argiento and De Iorio (2022). The proposed method is a finite analog of the inverse Gaussian processes proposed by Lijoi et al. (2005). We illustrate the performance of the proposed method for clustering, density estimation, and community detection, compared to existing MFM models based on Dirichlet distribution (e.g., Miller and Harrison, 2018; Geng et al., 2019; Argiento and De Iorio, 2022). The proposed method is robust against the choice of hyper-parameter $\alpha$, and provided reasonable estimates of the number of components and communities compared to the MFM based on the Dirichlet prior distribution. Moreover, the proposed method also has a reasonable predictive performance in the sense of density estimation by suppressing the appearance of empty components.
The drawbacks of the proposed method are as follows. Some parameters involved in the model do not have closed-form marginal distributions because it is not easy to marginalize out $U_{n}$. For example, when we focus on clustering, obtaining the interpretable prior distribution for the number of clusters is very important (see e.g., Zito et al., 2023) to incorporate subjective prior information. However, the proposed model cannot lead to a tractable marginal prior distribution for the number of clusters. In the mixture of finite mixtures, the model consists of a distribution over a simplex based on the normalization of independent random variables. Therefore, the correlation between categories cannot be properly modeled. The normalized inverse Gaussian distribution has a negative covariance as well as the Dirichlet distribution. It may be inappropriate for data with positive correlations between categories, such as the proportion of symbiotic organisms present, disease complication data, or gene expression data. To address such a problem, it may be necessary to construct the MFM in a more general framework that removes the assumption of independence in the Norm-IFPP by Argiento and De Iorio (2022). The construction of MFM models using a more flexible distribution over a simplex that can also express positive correlations such as the generalized Dirichlet distribution (Wong, 1998) is an interesting future topic. Furthermore, the proposed model can be applied to spatial data (e.g., Geng et al., 2021) and functional data (e.g., Hu et al., 2023). For network data, it is expected to extend the MFM to network with weighted edges, degree-corrected stochastic block models, and mixed membership stochastic block models.
Acknowledgement
This work was supported by Japan Society for the Promotion of Science, the establishment of university fellowships towards the creation of science technology innovation Grant Number JPMJFS2129. This work is partially supported by the Japan Society for the Promotion of Science (grant number: 21K13835).

### v2

@@ B.2 Additional Numerical Experiment
MFM-IGau
$M=1$ $M=2$ $M=3$ $M=4$ $M=5$ $M=6$ $M=7$ $M=8$ $M\geq 9$
$\alpha=10^{-1}$ $0.000$ $0.000$ $\bm{0.971}$ $0.024$ $0.004$ $0.001$ $0.000$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.059)$ $(0.046)$ $(0.021)$ $(0.003)$ $(0.000)$ $(0.000)$ $(0.000)$
$\alpha=10^{-2}$ $0.000$ $0.000$ $\bm{0.959}$ $0.036$ $0.004$ $0.001$ $0.000$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.063)$ $(0.054)$ $(0.008)$ $(0.004)$ $(0.002)$ $(0.001)$ $(0.000)$
$\alpha=10^{-3}$ $0.000$ $0.000$ $\bm{0.941}$ $0.048$ $0.007$ $0.002$ $0.001$ $0.000$ $0.001$
$(0.000)$ $(0.000)$ $(0.111)$ $(0.096)$ $(0.014)$ $(0.006)$ $(0.003)$ $(0.002)$ $(0.003)$
@@
$\gamma=10^{-1}$ $0.000$ $0.000$ $\bm{0.331}$ $0.301$ $0.187$ $0.098$ $0.047$ $0.021$ $0.016$
$(0.000)$ $(0.000)$ $(0.100)$ $(0.021)$ $(0.035)$ $(0.028)$ $(0.017)$ $(0.009)$ $(0.009)$
$\gamma=10^{-2}$ $0.000$ $0.000$ $0.098$ $0.172$ $\bm{0.187}$ $0.165$ $0.128$ $0.091$ $0.159$
$(0.000)$ $(0.000)$ $(0.021)$ $(0.024)$ $(0.013)$ $(0.005)$ $(0.009)$ $(0.012)$ $(0.037)$
$\gamma=10^{-3}$ $0.000$ $0.000$ $0.066$ $0.131$ $0.161$ $0.159$ $0.136$ $0.107$ $\bm{0.240}$
$(0.000)$ $(0.000)$ $(0.003)$ $(0.004)$ $(0.004)$ $(0.004)$ $(0.003)$ $(0.003)$ $(0.009)$
$\gamma=0.490$ $0.000$ $0.000$ $\bm{0.798}$ $0.170$ $0.028$ $0.004$ $0.001$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.070)$ $(0.048)$ $(0.019)$ $(0.005)$ $(0.001)$ $(0.000)$ $(0.000)$
$\gamma=0.352$ $0.000$ $0.000$ $\bm{0.717}$ $0.221$ $0.050$ $0.010$ $0.002$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.040)$ $(0.024)$ $(0.012)$ $(0.004)$ $(0.001)$ $(0.000)$ $(0.000)$
$\gamma=0.335$ $0.000$ $0.000$ $\bm{0.693}$ $0.237$ $0.056$ $0.012$ $0.002$ $0.000$ $0.000$
$(0.000)$ $(0.000)$ $(0.053)$ $(0.030)$ $(0.016)$ $(0.005)$ $(0.001)$ $(0.000)$ $(0.000)$